### v2

# COHESION: Composite Graph Convolutional Network with Dual-Stage Fusion for Multimodal Recommendation

Abstract.
Recent works in multimodal recommendations, which leverage diverse modal information to address data sparsity and enhance recommendation accuracy, have garnered considerable interest. Two key processes in multimodal recommendations are modality fusion and representation learning. Previous approaches in modality fusion often employ simplistic attentive or pre-defined strategies at early or late stages, failing to effectively handle irrelevant information among modalities. In representation learning, prior research has constructed heterogeneous and homogeneous graph structures encapsulating user-item, user-user, and item-item relationships to better capture user interests and item profiles. Modality fusion and representation learning were considered as two independent processes in previous work. This paper reveals that these two processes are complementary and can support each other. Specifically, powerful representation learning enhances modality fusion, while effective fusion improves representation quality. Stemming from these two processes, we introduce a COmposite grapH convolutional nEtwork with dual-stage fuSION for the multimodal recommendation, named COHESION. Specifically, it introduces a dual-stage fusion strategy to reduce the impact of irrelevant information, refining all modalities using behavior modality in the early stage and fusing their representations at the late stage. It also proposes a composite graph convolutional network that utilizes user-item, user-user, and item-item graphs to extract heterogeneous and homogeneous latent relationships within users and items. Besides, it introduces a novel adaptive optimization to ensure balanced and reasonable representations across modalities. Extensive experiments on three public datasets demonstrate the significant superiority of COHESION over various competitive baselines.
Multimodal, Recommender System, Dual-Stage Fusion
This work was supported by the Hong Kong UGC General Research Fund no. 17203320 and 17209822, and the project grants from the HKU-SCF FinTech Academy.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729927isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systems;
1. Introduction
The rapid development of the Internet has led to information explosion, making recommender systems an indispensable tool in human society. Traditional recommender systems rely on modeling user preferences through historical user-item interactions (Koren et al., 2021; Xu et al., 2024c; Chen et al., 2025; Xu et al., 2024a). However, the data sparsity problem always frustrates the accuracy of recommendations. As the types of information on social media become increasingly diverse, multimodal information has recently been used to alleviate the data sparsity problem (Lei et al., 2023; Mu et al., 2022; Xu et al., 2024d). A line of work (He and McAuley, 2016; Chen et al., 2017) integrates multimodal information as side information to enhance item representations. To better learn the representation, many works (Wang et al., 2021; Wei et al., 2019; Xu et al., 2024b) construct the user-item bipartite graph and utilize the graph convolution network (GCN) to enhance the representation learning in multimodal recommendations. Despite some previous works achieving notable recommendation performance via constructing the heterogeneous user-item graph, the relationships between user-user and item-item remain unexplored. Recently, DualGNN (Wang et al., 2021) constructs the user-user graph to explore the hidden preference pattern by similar users. LATTICE (Zhang et al., 2021) builds the item-item graph to capture semantically correlated signals. Both of these two homogeneous graphs improve the performance of recommendations. In this paper, we pose two comprehensive questions: Would exploiting both user-user and item-item graphs achieve even better performance? And how? To answer this question, we select the most standard GCN-based multimodal recommendation model MMGCN [^1] MMGCN is the earliest and simplest graph-based multimodal recommendation model, which only contains GCNs without extra components. (Wei et al., 2019) and carefully design its three variants [^2] All of the user-user graph, item-item graph, early fusion strategy, and late fusion strategy in this investigation are designed following a recent survey paper (Zhou et al., 2023c).: MMGCNu, MMGCNi, and MMGCNui, which incorporate only user-user, only item-item, and both user-user and item-item graphs, respectively. To assess the impact of fusion strategies, we implement the early fusion and late fusion strategies for each MMGCN variant, namely MMGCNe and MMGCNl, respectively. Table 1 shows the recommendation performance of eight MMGCN variants on the Baby dataset. In Table 1, our findings include: 1) The late fusion strategy outperforms the early fusion strategy when integrating only user-user or item-item graphs. Conversely, the early fusion strategy is more effective than late fusion when both graph types are integrated, possibly because late fusion permits deeper modality-specific learning, but it is more susceptible to the adverse effects of irrelevant inter-modality information. 2) Among the early fusion strategies, MMGCNui shows performance superior to MMGCNu and MMGCNi, suggesting that combining both user and item graphs facilitates more effective representation learning. These observations suggest that suboptimal fusion strategies in prior studies may have constrained the representation learning capabilities of the otherwise well-designed models. This hypothesis is empirically supported by the results presented in Table 6.
To liberate the representation learning ability of composite graph convolutional network, we introduce a COmposite grapH convolutional nEtwork with dual-stage fuSION for a multimodal recommendation, named COHESION. Specifically, we propose a novel dual-stage fusion strategy to mitigate the negative effects of irrelevant information among different modalities by refining all modalities using behavior modality in the early stage and fusing all modalities’ representations in the late stage. We believe that refining all modalities using behavior modality can enable each modality to map to a better semantic space. This is because the goal of multimodal recommendation is to predict user preferences instead of understanding multimodal content. Moreover, we propose the composite graph convolutional network to learn the representations for all modalities, including homogeneous and heterogeneous graphs. More specifically, we utilize a user-item graph to mine the relations between users and items and utilize user-user and item-item graphs to extract latent relations within users and items, respectively. Thanks to our dual-stage fusion strategy, our composite graph convolutional network liberates outstanding representation learning ability. Finally, we fuse all modality representations in the late stage and propose an adaptive optimization, which can make the representations learned from different modalities balanced. Extensive experiments on three widely used datasets demonstrate the superiority of COHESION over various competitive baselines.
Table 1. Performance of MMGCN variants in Baby dataset.
Metrics Recall@10 NDCG@10 Recall@20 NDCG@20
MMGCNe 0.0351 0.0589 0.0192 0.0249
[…]
MMGCNl 0.0378 0.0615 0.0200 0.0261
[…]
2. Related work
2.1. Multimodal Recommendation
Multimodal recommendation systems endeavor to harness multimodal information to enhance the accuracy of recommendations and to address the issue of data sparsity. Prior works (He and McAuley, 2016; Chen et al., 2017) alleviate the data sparsity challenge by extracting and utilizing visual content to enrich the representation of items based on matrix factorization techniques. Tran et al. (Tran and Lauw, 2022) extend this approach by incorporating textual content to further enrich item representations. More recently, several studies (Wang et al., 2021; Zhou and Shen, 2023) integrate information from multiple modalities concurrently. Drawing inspiration from traditional recommendation systems, Wei et al. (Wei et al., 2019, 2020) enhance the quality of user-item interactions through the implementation of a bipartite graph structure. To more effectively extract and analyze relationships between users and between items, DualGNN (Wang et al., 2021) introduces a user-user homogeneous graph to uncover latent relationships among users. Similarly, FREEDOM (Zhou and Shen, 2023) and LATTICE (Zhang et al., 2021) incorporate item-item homogeneous graphs to augment modality-specific information. MONET (Kim et al., 2024) and MARIO (Kim et al., 2022) leverage modality-aware attention and tailored GCN to explore items’ multimodal features. In the realm of exploring advanced structural possibilities, LGMRec (Guo et al., 2024) and DiffMM (Jiang et al., 2024) investigate the potential of hyper-graph structures and diffusion models, respectively, to significantly enhance the effectiveness of multimodal recommendation systems. To better alleviate the data sparsity problem, numerous studies (Tao et al., 2022; Zhou et al., 2023b; Xu et al., 2024d) integrate self-supervised learning tasks within the multimodal recommendation framework to discern hidden preference patterns across different modalities. However, it is important to note that self-supervised learning can also introduce irrelevant noisy information, which may include augmented self-supervised signals derived from user misclick behaviors or popularity biases, potentially reducing the model accuracy.
2.2. Multimodal Fusion
Multimodal information can effectively alleviate the data sparsity problem, and improve the accuracy of recommendations. However, multimodal information may lead to modal noise due to the feature of different modalities existing in different semantic spaces. Therefore, multimodal fusion plays a crucial role in noise mitigation. According to the timing of the modality fusion, most existing methods could be divided into two types: early fusion and late fusion. The early fusion methods (He and McAuley, 2016; Zhang et al., 2021) fuse the modality features with ID embedding before constructing the heterogeneous graph. However, the modality-specific features will be lost during training. Differently, the late fusion methods (Wei et al., 2019; Wang et al., 2021; Zhou and Shen, 2023) learn the representation of each modality, and fuse all the predicted ratings in the end.
With the rapid development of representation learning ability in recent multimodal recommendation models, predefined fusion strategies are more fragile to the negative impact of irrelevant information among modalities and are difficult to fuse different modalities effectively. To this end, we propose a tailored dual-stage fusion strategy, which effectively liberates the representation learning ability for multimodal recommendation models.
3. Methodology
Figure 1. The overview structure of COHESION.
In this section, we present our proposed COHESION [^1] The code will be released at: https://github.com/Jinfeng-Xu/COHESION after the SIGIR’25 conference. in detail. The overall architecture is shown in Fig. 1.
3.1. Preliminary
Let a set of users $u\in\mathcal{U}$ and a set of items $i$ = $\{i_{m}|i_{i},i_{t},i_{v}\}\in\mathcal{I}$, where $m\in\mathcal{M}$ is the modality, $\mathcal{M}$ is the set of modalities, and $i_{i}$, $i_{t}$, and $i_{v}$ represent the behavior, textual, and visual modality features, respectively. $\mathcal{G}$ = $(\mathcal{V},\mathcal{E})$ be a given graph with a node set $\mathcal{V}$ and an edge set $\mathcal{E}$, where $|\mathcal{V}|$ = $|\mathcal{U}|$ + $|\mathcal{I}|$. The user-item interaction matrix is denoted as $\mathcal{R}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|}$. Specifically, each entry $\mathcal{R}_{u,i}$ indicates whether the user $u$ is connected to item $i$, with a value of 1 representing a connection and 0 otherwise. The graph structure of $\mathcal{G}$ can be denoted as the adjacency matrix $\mathcal{A}\in\mathbb{R}^{(|\mathcal{U}|+|\mathcal{I}|)\times(|\mathcal{U}|+|\mathcal{I}|)}$:
(1) $\mathcal{A}=\left[\begin{array}[]{cc}0^{|\mathcal{U}|\times|\mathcal{U}|}&\mathcal{R}\\ \mathcal{R}^{T}&0^{|\mathcal{I}|\times|\mathcal{I}|}\end{array}\right].$
The symmetrically normalized matrix is $\mathcal{\tilde{A}}$ $=$ $\mathcal{D}^{-\frac{1}{2}}\mathcal{A}\mathcal{D}^{-\frac{1}{2}}$, where $\mathcal{D}$ represents a diagonal degree matrix. It is worth noting that we only consider three types of modalities in this paper, including behavior, textual, and visual modalities. However, we note it can be extended to more modalities easily.
3.2. Early Fusion
Multimodal information can effectively alleviate the data-sparsity problem for traditional recommendations. However, it also inevitably contains irrelevant information among modalities (Xv et al., 2024; Tang et al., 2019). To this end, we propose an early fusion strategy, which refines all modalities by behavior modality before constructing heterogeneous graphs. Our early fusion strategy can effectively build a bridge to mitigate the gap between different modalities by behavior modality. In particular, we first transform the feature embeddings $E_{i_{m}}$ = $\{E_{i_{i}},E_{i_{t}},E_{i_{v}}\}$ into the same dimension:
(2) $\dagger{E}_{i_{m}}=W_{m}^{\rhd}\sigma(W_{m}^{\lhd}E_{i_{m}}+b_{1})+b_{2},$
where $W_{m}^{\lhd}\in\mathbb{R}^{d_{m}\times 4d_{i}}$, $W_{m}^{\rhd}\in\mathbb{R}^{4d_{i}\times d_{i}}$, $b_{1}$, and $b_{2}$ are trainable parameters. $d_{m}$ and $d_{i}$ denote embedding dimensionality of modality $m$ and behavior modality $i$, respectively. $\sigma(\cdot)$ denotes the Leaky-ReLU function (Xu et al., 2020). Then we refine each modality embedding $\dagger{E_{i_{m}}}$ by behavior modality $\dagger{E_{i_{i}}}$ respectively:
(3) $\dagger{E}_{i_{m}^{i}}=\sqrt{|0.5\times((\dagger{E}_{i_{m}})^{2}+(\dagger{E}_{i_{i}})^{2})+\epsilon|},$
where $\dagger{E}_{i_{m}^{i}}$ denotes refined modality embedding for modality $m$, and $\epsilon$ is a positive infinitesimal to prevent zero vector. Through this tailored yet simple early fusion strategy, we enhance the representation learning capability of multimodal recommendation models by refining different modalities via the behavior modality. This early fusion strategy mitigates the gap between modalities, leading to more accurate and robust recommendations.
3.3. Heterogeneous Graph
To better learn the modality-specific user and item representations, we construct a user-item graph for each modality, which is denoted as $\mathcal{G}_{m}$ = $\{\mathcal{G}_{i},\mathcal{G}_{t},\mathcal{G}_{v}\}$. Following previous works (Wei et al., 2019; Zhou and Shen, 2023), we maintain the same graph structure $\mathcal{G}$ for different $\mathcal{G}_{m}$, but only retain the node features associated with a specific modality $m$. To alleviate the over-smoothing problem, we design a residual-based GCN for each user-item graph. The message propagation stage at the $l$-th graph convolution layer can be formulated as:
(4) $\dagger{E}^{(l)}_{m}=(\operatorname{Sim}(\mathcal{\tilde{A}}{\dagger{E}^{(l-1)}_{m}},\dagger{E}^{(0)}_{m})+\epsilon)\times\mathcal{\tilde{A}}{\dagger{E}^{(l-1)}_{m}},$
where $\dagger{E}^{(l)}_{m}$ denotes the refined representations of users and items in $l$-th graph convolution layer for the modality $m$, and $\dagger{E}^{(0)}_{m}$ $=$ $\operatorname{Con}(E_{u_{m}},\dagger{E}_{i_{m}^{i}})$ is the initial layer embedding, which concatenated by random initialed user embedding $E_{u_{m}}$ and refined modality embedding $\dagger{E}_{i_{m}^{i}}$. $\mathcal{\tilde{A}}$ is a symmetrically normalized matrix, $\operatorname{Sim}(\cdot)$ denotes the cosine similarity function, and $\epsilon$ is a small positive bias to avoid a zero vector. The final embedding for each modality is updated by summation:
(5) $\dagger{\bar{E}}_{m}=\sum_{l=0}^{L}\dagger{E}^{(l)}_{m},$
where $L$ is the total number of layers.
3.4. Late Fusion
In Section 1, we explored the impact of fusion strategies on representation learning. Previous work (Zhou et al., 2023d) also pointed out that existing fusion strategies may fail to capture modality-specific features and even corrupt the learned single-modality representation. To improve the quality of the modality fusion, we propose a tailored early fusion strategy to refine all modalities via behavior modality, which effectively alleviates the gap between modalities. After utilizing heterogeneous graphs to extract the hidden relationships between users and items, we fuse all the refined modality representations using attentive weights. Notably, our work is the first to effectively combine both early fusion and late fusion strategies to enhance the models’ representation learning capabilities.
(6) $\dagger{E}_{f}=\operatorname{Con}(\alpha_{m}\times\dagger{\bar{E}}_{m}|m=i,t,v),$
where $\dagger{E}_{f}$ is the final fused representation for all refined modalities. $\operatorname{Con}(\cdot)$ denotes the adaptive attention concatenation, and $\alpha_{m}$ is a learnable attention weight for each modality.
Conducting late fusion before enhancing the representations of users and items by homogeneous graphs can avoid the $|\mathcal{M}|$ times computation and storage cost of constructing homogeneous graphs for each modality separately and allow homogeneous graphs to capture features with modal synergy [^4] Partial information decomposition (Williams and Beer, 2010; Bertschinger et al., 2014) states the multivariate mutual information is decomposed into three forms of interaction (Uniqueness, Redundancy, and Synergy). Recent study (Dufumier et al., 2024) finds that interaction across different modalities allows homogeneous graphs to capture features with modal synergy. (Zhou and Shen, 2023; Dufumier et al., 2024).
3.5. Homogeneous Graph
To improve the recommendation performance by enhancing the representations for users and items, we construct the user-user and item-item homogeneous graphs.
3.5.1. User-User Graph
We construct the user-user graph to explore relations between users. Considering the computation overhead, we only capture a subset of users with the most similar preferences. To this end, the pairwise cosine similarity between all users is calculated as:
(7) $s_{u,\tilde{u}}=\frac{(e_{u})^{\top}e_{\tilde{u}}}{\|e_{u}\|\|e_{\tilde{u}}\|}.$
For each user $u$, we retain the edges with top-$k$ users $\mathcal{S}_{u}^{k}$ = $\{s_{u,\check{u}}|u,\check{u}\in\mathcal{U}\}$ among all user-user pairs:
(8) $s_{u,\check{u}}=\begin{cases}s_{u,\tilde{u}}&\text{ if }s_{u,\tilde{u}}\in\text{ top-}k(s_{u,v}|v\in\mathcal{U})\\ 0&\text{ otherwise }\end{cases},$
where $s_{u,v}|v\in\mathcal{U}$ represents the neighbor scores of $v$ for the user $u$. For the constructed user-user graph $\mathcal{G}^{u}$ = $\{\mathcal{U},\mathcal{S}_{u}^{k}\}$, we apply the Softmax function for propagation:
(9) $A^{(l)}_{u}=A^{(l-1)}_{u}+\sum_{s_{u,\check{u}}\in\mathcal{S}_{u}^{k}}\frac{\exp\left(s_{u,\check{u}}\right)}{\sum_{s_{u,\tilde{u}}\in\mathcal{S}_{u}}\exp\left(s_{u,\tilde{u}}\right)}A_{\tilde{u}}^{(l-1)},$
where $A^{(l)}_{u}$ is $l$-th layer representation of user $u$ learned from $\mathcal{G}^{u}$. $A^{(0)}_{u}$ is the user side $E_{f}(U)$ of $\dagger{E}_{f}$, which is obtained by $\dagger{E}_{f}[:|\mathcal{U}|]$.
3.5.2. Item-Item Graph
To explore relations between items, we construct an item-item graph to capture semantically correlative signals. However, the dense graph introduces huge noise, and the sparse graph leads to modality features under-exploited (Wang et al., 2021). Therefore, we only consider the top-$k$ most similar items $\mathcal{S}_{i}^{k}$ = $\{s_{i,\check{i}}|i,\check{i}\in\mathcal{I}\}$ for each item $i\in\mathcal{I}$. To this end, the pairwise cosine similarity between all items is calculated as follows:
(10) $s_{i,\tilde{i}}=\frac{(s_{i})^{\top}s_{\tilde{i}}}{\|s_{i}\|\|s_{\tilde{i}}\|}.$
Same as the user-user graph, for each item $i$, we retain the edges with top-$k$ items $\mathcal{S}_{i}^{k}$ = $\{s_{i,\check{i}}|i,\check{i}\in\mathcal{I}\}$ among all item-item pairs:
(11) $s_{i,\check{i}}=\begin{cases}s_{i,\tilde{i}}&\text{ if }s_{i,\tilde{i}}\in\text{ top-}k(s_{i,j}|j\in\mathcal{I})\\ 0&\text{ otherwise }\end{cases},$
where $s_{i,j}|j\in\mathcal{I}$ represents the neighbor scores for the item $i$. We further construct a item semantic graph $\mathcal{G}^{i}$ = $\{\mathcal{I},\mathcal{S}_{i}^{k}\}$, and apply weighted aggregation function for propagation:
(12) $A^{(l)}_{i}=\sum_{s_{i,\check{i}}\in S_{i}^{k}}s_{i,\check{i}}\times A_{\check{i}}^{(l-1)},$
where $A^{(l)}_{i}$ is $l$-th layer representation of item $i$ learned from $\mathcal{G}^{i}$. $A^{(0)}_{i}$ is the item side $E_{f}(I)$ of $\dagger{E}_{f}$, which can be obtained by $\dagger{E}_{f}[|\mathcal{U}|:]$. Then, we enhance the user and item representations by user-user and item-item graphs, respectively:
(13) $E_{f}(U)=\dagger{E}_{f}(U)+A_{u}^{L_{u}},\quad E_{f}(I)=\dagger{E}_{f}(I)+A_{i}^{L_{i}},$
(14) $E_{f}=\operatorname{Con}(E_{f}(U)|E_{f}(I)),$
where $A_{u}^{L_{u}}$ and $A_{i}^{L_{i}}$ are the final layer results for the user-user graph and the item-item graph, respectively. $\operatorname{Con}(\cdot)$ denotes a concatenation operation. $E_{f}(U)$ denotes the final user representation, which can be obtained by $\dagger{E}_{f}[:|\mathcal{U}|]$, and $E_{f}(I)$ denotes the final item representation, which can be obtained by $\dagger{E}_{f}[|\mathcal{U}|:]$.
3.6. Adaptive Optimization
Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2012) are widely adopted to optimize the trainable parameters. In particular, BPR increases the gap between user $u$’s predicted ratings of the positive item $p$ and negative item $n$ for each triplet $(u,p,n)\in\mathcal{D}$, where $\mathcal{D}$ denotes the training set. Positive item $p$ is the item that has interaction with user $u$, and negative item $n$ is the item randomly sampled from items without any interaction with user $u$.
Previous work (Dong et al., 2024) pointed out that the BPR loss is limited to optimizing the overall user preferences and does not directly optimize the learning of modality-specific preferences, i.e., user interests specific to each modality. As a result, there could be some modality-specific interest that is not well-learned, potentially leading to unreliable final user rating predictions. To better learn the features of all modalities, we propose an adaptive BPR loss function, which pays more attention to the weak modality. The adaptive BPR loss function is defined as follows:
(15) $w_{u,p,n}^{m}=1-\frac{\exp(y_{u,p}^{m}-y_{u,n}^{m})}{\sum_{m\in M}\exp(y_{u,p}^{m}-y_{u,n}^{m})},$
(16) $\mathcal{L}=-\sum_{(u,p,n)\in\mathcal{D}}\log(\sigma(\sum_{m\in M}{w_{u,p,n}^{m}(y_{u,p}^{m}-y_{u,n}^{m})}))+\lambda(\|\mathbf{\Theta}\|_{2}^{2}),$
where $w_{u,p,n}^{m}$ is the weight for each modality calculated by reversing the rating gap, and the rating gap is the distance between user $u$’s predicted ratings of the positive item $y_{u,p}^{m}$ and the negative item $y_{u,n}^{m}$ for each modality. $\sigma(\cdot)$ is the Sigmoid function, $\lambda$ is a hyper-parameter for regularization term, and $\mathbf{\Theta}$ denotes model parameters.
To provide a clearer overview of our COHESION, we summarize the learning process of COHESION in Algorithm 1.
Algorithm 1 Learning Process of COHESION
1:  Input: $\mathcal{U}$, $\mathcal{I}$, $\mathcal{M}$, $\mathcal{G}$, layer number $L$ of heterogeneous graph $\mathcal{G}$, layer number $L_{u}$ of user-user graph $\mathcal{G}^{u}$, and layer number $L_{i}$ of item-item graph $\mathcal{G}^{i}$
2:  Output: Optimization loss $\mathcal{L}$
3:  Initialize $E_{u}$, $E_{i_{m}}$;
4:  Transform item embeddings of all modalities into the same dimension $\dagger{E}_{i_{m}}$leftarrow $E_{i_{m}}$ with Eq.2;
5:  (Early Fusion) Refine item embeddings of each modality by behavior modality $\dagger{E}_{i_{m}^{i}}$leftarrow $\dagger{E}_{i_{m}}$ with Eq.3;
6:  for $l=1...L$ do
7:     Conduct message propagation within the residual-based heterogeneous graph $\dagger{E}^{(l)}_{m}$leftarrow $\dagger{E}^{(l-1)}_{m}$ with Eq.4;
8:  end for
9:  Get embedding $\dagger{\bar{E}}_{m}$ for each modality with Eq.5;
10:  (Late Fusion) Attentively fuse all modality embeddings $\dagger{E}_{f}$leftarrow $\dagger{\bar{E}}_{m}$ with Eq.6;
11:  Construct the user-user graph $\mathcal{G}^{u}$ with Eq.7 and Eq.8.
12:  Construct the item-item graph $\mathcal{G}^{i}$ with Eq.10 and Eq.11.
13:  for $l=1...L_{u}$ do
14:     Conduct message propagation within the user-user homogeneous graph $A^{(l)}_{u}$leftarrow $A^{(l-1)}_{u}$ with Eq.9;
15:  end for
16:  for $l=1...L_{i}$ do
17:     Conduct message propagation within the item-item homogeneous graph $A^{(l)}_{i}$leftarrow $A^{(l-1)}_{i}$ with Eq.12;
18:  end for
19:  Get final enhanced embedding $E_{f}$ by both user-user and item-item graphs with Eq.13 and Eq.14.
20:  Calculate adaptive BPR loss $\mathcal{L}$ with Eq.15 and Eq.16.
4. Experiments
In this section, we conduct extensive experiments on some widely used real-world datasets. Experiment results can evaluate the following questions: RQ1: How does COHESION perform compared with various state-of-the-art models? RQ2: How do the key components in COHESION influence the performance of recommendation accuracy? RQ3: Can COHESION effectively mitigate the irrelevant information among modalities? RQ4: How does COHESION perform compared with various state-of-the-art models in different sparse data scenarios? RQ5: How efficient is COHESION compared with various state-of-the-art models? RQ6: How sensitive is COHESION under the perturbation of hyper-parameters?
Table 2. Statistics of three experimented datasets with modalities item Visual(V) and Textual(T).
Dataset Baby Sports Clothing
Modality V T V T V T
Embed Dim 4,096 384 4,096 384 4,096 384
User 19,445 35,598 39,387
Item 7,050 18,357 23,033
Interaction 160,792 296,337 278,677
Sparsity 99.88% 99.95% 99.97%
4.1. Settings
4.1.1. Datasets
Our experiments are conducted on three real-world datasets: Baby, Sports, and Clothing from the Amazon dataset (McAuley et al., 2015), which include visual and textual modalities for each item. Table 2 shows the statistics of these datasets. Following most previous works (Yi and Chen, 2021; Zhou et al., 2023b; Xu et al., 2025), we adopt the 5-core setting to filter users and items for each dataset. We use the pre-trained sentence-transformer to extract textual features with 384 dimensions and use the published 4096-dimensional visual features as in MMRec (Zhou, 2023). For each dataset, we use the ratio 8:1:1 to split the historical interactions randomly for training, validation, and testing.
Table 3. Performance comparison of baselines and COHESION in terms of Recall@K (R@K) and NDCG@K (N@K). The superscript indicates the improvement is statistically significant where the p-value is less than 0.01.
Datasets Baby Sports Clothing
Model R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20
MF-BPR 0.0357 0.0575 0.0192 0.0249 0.0432 0.0653 0.0241 0.0298 0.0187 0.0279 0.0103 0.0126
LightGCN 0.0479 0.0754 0.0257 0.0328 0.0569 0.0864 0.0311 0.0387 0.0340 0.0526 0.0188 0.0236
SimGCL 0.0513 0.0804 0.0273 0.0350 0.0601 0.0919 0.0327 0.0414 0.0356 0.0549 0.0195 0.0244
LayerGCN 0.0529 0.0820 0.0281 0.0355 0.0594 0.0916 0.0323 0.0406 0.0371 0.0566 0.0200 0.0247
VBPR 0.0423 0.0663 0.0223 0.0284 0.0558 0.0856 0.0307 0.0384 0.0281 0.0415 0.0158 0.0192
MMGCN 0.0378 0.0615 0.0200 0.0261 0.0370 0.0605 0.0193 0.0254 0.0218 0.0345 0.0110 0.0142
DualGNN 0.0448 0.0716 0.0240 0.0309 0.0568 0.0859 0.0310 0.0385 0.0454 0.0683 0.0241 0.0299
SLMRec 0.0529 0.0775 0.0290 0.0353 0.0663 0.0990 0.0365 0.0450 0.0452 0.0675 0.0247 0.0303
LATTICE 0.0547 0.0850 0.0292 0.0370 0.0620 0.0953 0.0335 0.0421 0.0492 0.0733 0.0268 0.0330
FREEDOM 0.0627 0.0992 0.0330 0.0424 0.0717 0.1089 0.0385 0.0481 0.0628 0.0941 0.0341 0.0420
BM3 0.0564 0.0883 0.0301 0.0383 0.0656 0.0980 0.0355 0.0438 0.0422 0.0621 0.0231 0.0281
LGMRec 0.0639 0.0989 0.0337 0.0430 0.0719 0.1068 0.0387 0.0477 0.0555 0.0828 0.0302 0.0371
COHESION 0.0680 0.1052 0.0354 0.0454 0.0752 0.1137 0.0409 0.0503 0.0665 0.0983 0.0358 0.0438
Improv. 6.42% 6.05% 5.04% 5.58% 4.59% 4.41% 5.68% 4.57% 5.89% 4.46% 4.99% 4.28%
4.1.2. Baseline
To evaluate the effectiveness of our COHESION, we compare it with the following methods, which can be divided into two groups. We detailed all of them as follows:
1) Traditional recommendation models:
  • MF-BPR (Rendle et al., 2012) utilizes the BPR loss to optimize the classic collaborative filtering method, which learns user and item representations by the matrix factorization method.
  • LightGCN (He et al., 2020) excludes some unnecessary components from GCN-based collaborative filtering to make it more appropriate for recommendation.
  • SimGCL (Yu et al., 2022) proposes a graph contrastive learning method, which directly injects random noises into the representation.
  • LayerGCN (Zhou et al., 2023a) utilizes the residual connection to build a layer-refined GCN to alleviate LightGCN’s over-smoothing problem.
2) Multimodal recommendation models:
  • VBPR (He and McAuley, 2016) integrates the visual and textual features with ID embeddings as side information for each item, which can be seen as multimodal matrix factorization.
  • MMGCN (Wei et al., 2019) utilizes GCN for each modality to learn the modality-specific features, and then fuses all the user predicted ratings for all modalities to get the final predicted rating.
  • DualGNN (Wang et al., 2021) proposes a user-user graph to explore the hidden preference pattern.
  • LATTICE (Zhang et al., 2021) proposes an item-item graph to capture semantically correlated signals among items.
  • FREEDOM (Zhou and Shen, 2023) optimizes LATTICE by freezing the item-item graph and denoising the user-item graph.
  • SLMRec (Tao et al., 2022) proposes a self-supervised framework for the multimodal recommendation, which builds a node self-discrimination task to uncover the hidden multimodal patterns of items.
  • BM3 (Zhou et al., 2023b) simplifies the self-supervised framework by directly perturbing the representation through a dropout mechanism.
  • LGMRec (Guo et al., 2024) utilizes local embeddings with local topological information and global embeddings with hypergraphs.
4.1.3. Evaluation Metrics
For a fair comparison, we follow the settings of previous works (Wang et al., 2021; Zhou et al., 2023b; Zhou and Shen, 2023) to adopt two widely-used evaluation metrics for top-$K$ recommendation: Recall@K and NDCG@K. We report the average scores for all users in the test dataset under both K = 10 and K = 20, respectively.
4.1.4. Implementation Details
Following the existing work (Zhou, 2023), we fix the user and item embedding size to 64 for all models, and then initialize them with the Xavier method (Glorot and Bengio, 2010). Meanwhile, we optimize all models with Adam optimizer (Kingma and Ba, 2014). For each baselines, we perform a grid search for all the hyper-parameters following its published paper to find the optimal setting. All models are implemented by PyTorch and evaluated on an RTX 4070Ti GPU card. For our proposed COHESION, we perform a grid search on the learning rate in {1e-1, 1e-2, 1e-3, 1e-4}, the regularization weight $\lambda$ in {1e-1, 1e-2, 1e-3, 1e-4}, and residual-based GCN layers $L$ in {1, 2, 3, 4}. We empirically fix the number of GCN layers in the homogeneous graph (both user-user graph and item-item graph) with $L_{u}/L_{i}$ = 1 (Zhang et al., 2021; Wang et al., 2021). The $k$ of top-$k$ in the user-user graph is also empirically set as 10 (Zhou and Shen, 2023; Zhang et al., 2021; Wang et al., 2021). To avoid over-fitting, the early stopping and total epochs are fixed at 20 and 1000, respectively. Following MMRec (Zhou, 2023), we use Recall@20 on the validation dataset as the indicator for best record updating.
4.2. Performance Comparison (RQ1)
Table 3 shows the performance comparison of our proposed COHESION with other baselines on three widely used datasets. The optimal and optimal results are indicated in bold and underlined, respectively. We have the following observations:
  • Most multimodal models achieve better performance than traditional recommendation models, which verifies that multimodal information can effectively alleviate the data sparsity problem in traditional recommendation systems.
  • Some multimodal recommendation methods result in inconsistent performance across various scenarios. For example, LATTICE outperforms SLMRec on the Baby dataset, but is less effective on the Sports dataset. The self-supervised learning works better in the Sports dataset than in the Baby dataset.
  • COHESION makes a huge improvement over all multimodal recommendation models. This validates the effectiveness of our COHESION. Our dual-stage fusion strategy effectively bridges different modalities by refining each modality via behavior modality. Our well-designed composite graph convolutional network leverages both heterogeneous and homogeneous graphs to improve the learning representation ability. Moreover, we introduce an adaptive optimization, which is designed to enhance the comprehensive learning of under-learned modalities. As a result, COHESION outperforms existing multimodal methods. For example, COHESION improves the strongest baseline (FREEDOM or LGMRec) in terms of R@20 on the Baby and Sports datasets by 6.05% and 4.41%, respectively.
  • DualGNN and LATTICE both achieve competitive performance by adding a user-user and item-item graph, respectively. Our COHESION significantly outperforms these two methods. We owe our superiority to our novel dual-stage fusion strategy liberates the representational capabilities of a well-designed composite graph convolutional network.
4.3. Ablation Study (RQ2)
In this section, we empirically evaluate the effectiveness of different components of COHESION.
4.3.1. Effect of Early Fusion
To investigate the effects of the early fusion component in COHESION, we set up the following model variants:
  • COHESION-itv: Removing the entire early fusion process.
  • COHESION-tv: Removing refinement operation of textual and visual modalities in the early fusion process.
  • COHESION-iv: Removing the refinement operation of behavior and visual modalities in the early fusion process.
  • COHESION-it: Removing the refinement operation of behavior and textual modalities in the early fusion process.
  • COHESION-v: Removing the refinement operation of visual modality in the early fusion process.
  • COHESION-t: Removing the refinement operation of textual modality in the early fusion process.
  • COHESION-i: Removing the refinement operation of behavior modality in the early fusion process.
Figure 2. Effect of early fusion.
Fig. 2 shows the results (Recall@20) of these variants and COHESION on all three datasets. We have the following observations:
  • 1) Results of COHESION-itv $<$ COHESION-tv, COHESION-iv $<$ COHESION-v, and COHESION-it $<$ COHESION-t showing that refining behavior modality also leads to a slight improvement in recommendation performance.
  • 2) COHESION performs better than COHESION-i, COHESION-v, and COHESION-t, which shows that refining any modality can lead to improve the recommendation accuracy.
  • 3) COHESION outperforms all seven variants, verifying the effectiveness of our early fusion strategy.
Table 4. Effect of homogeneous graphs in terms of Recall@K (R@K) and NDCG@K (N@K).
Dataset Variants R@10 R@20 N@10 N@20
Baby COHESION 0.0680 0.1052 0.0354 0.0454
COHESION-u 0.0660 0.1024 0.0349 0.0443
COHESION-i 0.0564 0.0884 0.0307 0.0389
COHESION-ui 0.0547 0.0870 0.0291 0.0375
Sports COHESION 0.0752 0.1137 0.0409 0.0503
COHESION-u 0.0705 0.1086 0.0384 0.0482
COHESION-i 0.0622 0.0963 0.0337 0.0425
COHESION-ui 0.0603 0.0937 0.0327 0.0413
Clothing COHESION 0.0665 0.0983 0.0358 0.0438
COHESION-u 0.0630 0.0942 0.0339 0.0417
COHESION-i 0.0499 0.0744 0.0271 0.0331
COHESION-ui 0.0478 0.0701 0.0262 0.0319
4.3.2. Effect of Homogeneous Graphs
To identify the effectiveness of homogeneous graphs in COHESION, we set up the following variants: 1) COHESION-ui, which removes both user-user and item-item graphs. 2) COHESION-i, which removes only the item-item graph. 3) COHESION-u, which removes only the user-user graph. As illustrated in Table 4, we observe that both user-user and item-item graphs could improve the performance of recommendations. The item-item graph can better represent the user’s preference than the user-user graph, which is identified by the results COHESION-ui $<$ COHESION-i $<$ COHESION-u $<$ COHESION. Note that user-user and item-item graphs can improve recommendation accuracy together to achieve higher performance.
Table 5. Effect of adaptive optimization in terms of Recall@K (R@K) and NDCG@K (N@K).
Dataset Variants R@10 R@20 N@10 N@20
Baby DualGNN 0.0448 0.0716 0.0240 0.0309
DualGNN+ 0.0492 0.0753 0.0261 0.0328
LATTICE 0.0547 0.0850 0.0292 0.0370
LATTICE+ 0.0573 0.0881 0.0303 0.0388
Sports DualGNN 0.0568 0.0859 0.0310 0.0385
DualGNN+ 0.0587 0.0894 0.0316 0.0394
LATTICE 0.0620 0.0953 0.0335 0.0421
LATTICE+ 0.0634 0.0974 0.0343 0.0430
Clothing DualGNN 0.0454 0.0683 0.0241 0.0299
DualGNN+ 0.0499 0.0737 0.0270 0.0334
LATTICE 0.0492 0.0733 0.0268 0.0330
LATTICE+ 0.0524 0.0787 0.0285 0.0354
4.3.3. Effect of Adaptive Optimization
Our adaptive optimization could serve as a play-and-plug component for other models to improve their performance. To evaluate the effectiveness of adaptive optimization in other models, we design the corresponding model variants, named DualGNN+ and LATTICE+, where symbol + means using adaptive optimization. As shown in Table 5, our adaptive optimization achieves significant improvement for all models on all datasets.
Table 6. Performance comparison on all three datasets (Recall@20). The notation Visual means only adding visual modality, Textual means only adding textual modality, and Multimodal means adding both visual and textual modalities. Improv. denotes the improvement of adding multimodal compared to adding the optimal single modality.
Dataset Modality VBPR MMGCN DualGNN SLMRec LATTICE FREEDOM BM3 LGMRec COHESION
Baby Visual 0.0665 0.0634 0.0830 0.0789 0.0779 0.0816 0.0806 0.0954 0.0878
Textual 0.0638 0.0623 0.0943 0.0797 0.0880 0.0980 0.0842 0.0997 0.1002
Multimodal 0.0663 0.0615 0.0716 0.0775 0.0848 0.0992 0.0833 0.0989 0.1052
Improv. -0.30% -3.09% -31.17% -2.84% -3.77% +1.22% -1.08% -0.80% +4.99%
Sports Visual 0.0849 0.0574 0.0896 0.0928 0.0902 0.0879 0.0973 0.1059 0.1085
Textual 0.0714 0.0623 0.0958 0.1011 0.0978 0.1083 0.0989 0.1093 0.1099
Multimodal 0.0856 0.0605 0.0859 0.0990 0.0953 0.1089 0.0980 0.1068 0.1137
Improv. -0.82% -2.89% -10.33% -2.08% -2.56% +0.55% -0.91% -2.29% +3.46%
Clothing Visual 0.0433 0.0363 0.0730 0.0662 0.0705 0.0816 0.0608 0.0805 0.0928
Textual 0.0397 0.0358 0.0784 0.0698 0.0767 0.0937 0.0650 0.0847 0.0956
Multimodal 0.0415 0.0345 0.0683 0.0675 0.0733 0.0941 0.0621 0.0828 0.0983
Improv. -4.16% -4.96% -12.77% -3.30% -4.43% +0.43% -4.46% -2.24% +2.82%
4.4. Effectiveness of Mitigating Irrelevant Information among Modalities (RQ3)
The negative effect of irrelevant information among modalities will decrease modal representation learning ability and even lead to a weaker performance for incorporating multiple modalities than incorporating a single modality. Table 6 shows that most existing models suffer from this dilemma, we observe that only FREEDOM and our COHESION achieve higher performance for incorporating multiple modalities and that COHESION made a much more significant improvement of 4.99%. It verifies that our early fusion strategy bridges different modalities to mitigate the negative effect of irrelevant information among modalities.
Figure 3. Visualization of distribution through t-SNE. We randomly sample 500 items from Baby dataset. Red, green, and blue represent the ID, textual, and visual modalities.
To gain a better understanding of the advantages of our dual-stage strategy for mitigating irrelevant information among modalities, we take a step: randomly pick 500 items from the Baby dataset and use the t-SNE (Van der Maaten and Hinton, 2008) algorithm to map the representations of late fusion strategy and our dual-stage strategy into a two-dimensional space, respectively. Fig. 3 visualizes the distributions of all modalities. The distributions of all modalities from our dual-stage fusion strategy exhibit significantly greater similarity compared to those from the late fusion strategy, further validating the effectiveness of our dual-stage fusion approach in aligning modality representations.
Figure 4. Sparsity degree analysis on three datasets.
4.5. Different Sparse Data Scenarios (RQ4)
In this section, we verify the effectiveness of COHESION in different sparse data scenarios. We conduct experiments on sub-datasets with varying levels of data sparsity using all three datasets. We compare the performance of COHESION against three outstanding baselines (LATTICE, FREEDOM, LGMRec). We split user groups by the number of interactions in the training set (e.g., the first group consists of users with 0-5 interacted items). Fig. 4 shows that COHESION consistently outperforms all baselines on all datasets with different degrees of sparsity, which demonstrates the effectiveness in different sparse data scenarios.
Table 7. Comparison of COHESION against competitive baselines on efficiency. Parameter denotes the learnable parameter for the model, and Time denotes the total duration of each iteration, including both the training and evaluation phases.
Dataset Metrics VBPR MMGCN DualGNN SLMRec LATTICE FREEDOM BM3 LGMRec COHESION
Baby Time (s/epoch) 2.12 3.96 5.48 3.75 4.13 2.77 2.55 5.93 4.47
Parameter (M) 3.2 1.4 3.8 2.0 33.6 33.6 33.6 38.6 37.0
Sports Time (s/epoch) 3.55 11.01 15.90 6.14 11.90 5.86 5.39 8.98 7.91
Parameter (M) 6.0 1.4 5.9 3.8 86.0 86.0 86.0 99.9 91.5
Clothing Time (s/epoch) 3.90 13.04 16.93 6.94 14.10 6.29 6.09 10.02 9.05
Parameter (M) 6.8 1.4 6.4 4.3 107.5 107.5 107.5 126.1 113.4
4.6. Efficiency Analysis (RQ5)
4.6.1. Time Complexity
[…]
4.6.2. Computational Resource
Moreover, we report the training time (including both computation time and evaluating time) and model parameters for our COHESION and baselines in Table 7. We observe that our COHESION achieves a significant improvement over competitive baselines (LATTICE, FREEDOM, BM3, and LGMRec) while requiring comparable computation resources and training time. This observation underscores the efficiency of our COHESION model, highlighting its ability to achieve superior recommendation performance without necessitating an increase in resource consumption or training time, thereby demonstrating its practical applicability in multimodal recommendation scenarios.
4.6.3. Convergency
In Figure 5, we show the training curves of our COHESION and competitive baselines (LATTICE, FREEDOM, and LGMRec) on all three datasets, as the numbers of iterations and epochs increase. The faster convergence of our COHESION method is obviously observed, which suggests the advantage of COHESION in training efficiency while maintaining superior recommendation accuracy. This enhanced performance can be attributed to our innovative dual-stage fusion strategy, which effectively mitigates the influence of irrelevant information across modalities and brings a positive effect for model optimization to achieve fast convergence. Our dual-stage fusion strategy not only expedites the convergence speed but also enhances recommendation performance.
Figure 5. Convergence study in terms of Recall@20.
4.7. Hyper-parameter Sensitivity Study (RQ6)
4.7.1. Number of Layer $L$ of Heterogeneous Graph
As shown in Fig. LABEL:Layers, we find that the optimal number of layer $L$ for Baby, Sports, and Clothing datasets are 1, 2, and 2, respectively. The reason is that the degree of association between users and items is smaller in the Baby dataset than in the Sports and Clothing datasets. It is important to emphasize that for all other multimodal recommendation methods using GCN, the best number of layers of GCN is 1 for all these three datasets. It verifies that residual-based GCN can alleviate the over-smoothing problem to some extent.
4.7.2. The Learning Rate and Regularization Weight
Fig. LABEL:fig:5 shows the results of COHESION under different learning rates and regularization weights on Baby, Sports, and Clothing datasets. We can observe that 1e-3, 1e-4, and 1e-3 are the suggested learning rates for the Baby, Sports, and Clothing datasets, respectively. Moreover, the suggested regularization weights are 1e-4, 1e-3, and 1e-3 for the Baby, Sports, and Clothing datasets, respectively.
5. Conclusion
In this paper, we examine the limitations of the fusion strategy in previous works and further reveal the complementary capability between fusion strategy and representation learning. To this end, we propose a composite graph convolutional network with a dual-stage fusion strategy for multimodal recommendation, named COHESION. More specifically, our well-designed composite graph convolutional network achieves outstanding representation learning ability by leveraging both heterogeneous and homogeneous graphs. Moreover, our tailored dual-stage fusion strategy further liberates the representation learning ability of our composite graph convolutional network by mitigating the negative effect of irrelevant information among modalities. Finally, we propose an adaptive optimization to control the balance of attention among all modalities. We conduct comprehensive experiments and ablation studies to demonstrate the effectiveness of our COHESION.